- C R2 R1 + Vi(jw) Vo R3 H HI-Assuming an ideal op-amp (1)find the 3dB frequency of the circuit, if R1=1.5 KΩ, R2=13.2 KΩ, R3=20 KΩ, and C=5 nF

2)Consider the above circuit with the component values: R1=13.2 KΩ, R2= 795.77 Ω, R3=20 KΩ, and C=100 nF.  Find the angle of the transfer function in degrees at 2KHz .
Answer #1

a میں کر с th AI VN I + funt the Center into ) terminals of op-amp R2 RI Vo(s) to be ideal nsidering op - amp Can obseave tha

R2 of and The paralle will provide с Combination impedence Rallisce R2. Sc R2 + Sc R2 Z It SCR2 Now Vo (s) ZIT VN (Applying K

1+ W= Relai Vols) Vils) 1/R2C. at R2C Rec Vols) Relri Vics) yz which obtained frequency So cut off frequency at 3-dB is fc =

(2) Transfer function R2 12/R1 Vout (s) Vin(s) 1+ jwcR2 $ = - tan! (wcR2) = - tani ton (20 2UX QX103 x 400xio X 795.77) 450

Please comment if you have any doubts...

Please give me upvote if you like my answer...tq...